OpenStudy (anonymous):

It is known that the Maclaurin's series for f(x) is 1 + 3x + 5x² + 7x³ + ... . Find the value of f'(0), f''(0) and f'''(0).

OpenStudy (anonymous):

I differentiate 3 times and can get the values of f'(0) and f''(0) but not f'''(0). Why is this so ? :(

OpenStudy (skullpatrol):

Because that is what the formula says you should do :)

OpenStudy (skullpatrol):

The first time you differentiate f(x) you get f'(x), right @HatcrewS ?

zepdrix (zepdrix):

\[\Large\rm f(x)=1+3x+5x^2+7x^3+...\] \[\Large\rm f(x)=\sum_{0}^{\infty}\frac{f^n(0)}{n!}x^n\] Is this what you're having trouble with? The third derivative? Maybe I misunderstood your question though -_- \(\Large\rm \implies\quad \frac{f^{(3)}(0)}{3!}=7x^3\)

OpenStudy (joannablackwelder):

3=f'(0) using the idea that a1=f'(0) for functions where x exists at 0.

zepdrix (zepdrix):

Woops I left out the x^3 on the left lol.

OpenStudy (anonymous):

@zepdrix i differentiate til i get d³y/dx³ and sub in x = 0

zepdrix (zepdrix):

Oh you're not using the formula? :3 my bad.

OpenStudy (joannablackwelder):

And f''(0)= a2*2! which would be 5*2=10

OpenStudy (anonymous):

f'(x) = 3 + 10x + 21x² f''(x) = 10 + 42x f'''(x) = 42

zepdrix (zepdrix):

Looks good c:

OpenStudy (anonymous):

thank you :D

OpenStudy (skullpatrol):

f'''(x) is the constant 42 :)

OpenStudy (anonymous):

thanks everyone for helping! :DD

OpenStudy (skullpatrol):

Thanks for asking :)

ganeshie8 (ganeshie8):

Don't ignore the remaining terms : f'(x) = 3 + 10x + 21x² + ... f''(x) = 10 + 42x + ... f'''(x) = 42 + .... they vanish olny after plugging in x=0 (just nitpicking :P)

OpenStudy (skullpatrol):

Good point @ganeshie8